OpenStudy (karamarie02):

please answer as fast as possible!!!!!!will fan and medal solve. -14=2/3t what does t equal?

TheSmartOne (thesmartone):

what is the reciprocal of 2/3

TheSmartOne (thesmartone):

@karamarie02

TheSmartOne (thesmartone):

hint: the reciprocal of a/b is b/a

OpenStudy (karamarie02):

3/2 @TheSmartOne

TheSmartOne (thesmartone):

correct so we need to multiply both sides by the recirpocal of 2/3 which is 3/2 that is because we need to isolate 't' so we can get the answer

OpenStudy (karamarie02):

that's what I don't get @TheSmartOne

TheSmartOne (thesmartone):

\[-14 \times \frac{3}{2} = \frac{2}{3} \times \frac{3}{2} t\] We can simplify the right side to just 't' can you simplify the left side?